Three times a first number decreased by a second number is 3.the first number increases by twice the second number is 15.find the numbers
Answer by addingup(3677)

3x-y= 3
x+2y= 15 Let's multiply both sides by 3:
3x+6y= 45 Now subtract this equation from the first:
-----------------------------
3x-y= 3
- 3x-6y= -45 (watch your signs, you are subtracting so numbers turn negative)
___ -7y= -42
____ y= 6
So we have that y, the second number, is 6. Now back to the top, to the first equation:
--------------------------------
3x-y= 3 Using y= 6:
3x-6= 3
3x= 9
x= 3 This is the first number
----------------------------
Check:
3(3)-6= 3
9-6= 3 OK, we got this correct
3+2(6)= 15
3+12= 15
15=15 This one is also correct. We have the right answers.